4(2x-3)=-2(3x-8)
what is x?

Found 2 solutions by algebrahouse.com, Maths68:
Answer by algebrahouse.com(1659)

4(2x - 3) = -2(3x - 8)
8x - 12 = -6x + 16 {used distributive property}
14x - 12 = 16 {added 6x to both sides}
14x = 28 {added 12 to both sides}
x = 2 {divided both sides by 14}

For more help from me, visit: www.algebrahouse.com



Answer by Maths68(1474)

4(2x-3)=-2(3x-8)
8x-12=-6x+16
8x+6x=16+12
14x=28
14x/14=28/14
x=2
Check
4(2x-3)=-2(3x-8)
4(2(2)-3)=-2(3(2)x-8)
4(4-3)=-2(6-8)
4(1)=-2(-2)
4=4